# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

### v2

@@ 3.1 Hamilton-Pontryagin Principle and Lagrangian-Eulerian Formulation
and conclude that the coupled system (39) is equivalent to the system (1) with pressure $p$.
(II) We compute the partial derivatives $\partial_{\xi}H$, $\partial_{X}H$. Recall the definition of variational derivative:
$\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}H(\xi+\varepsilon\delta\xi,X,u)=\int_{D}[\frac{\partial H}{\partial\xi}(\xi,X,u)\cdot\delta\xi]\rho_{0}(a)da$
@@
$\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}H(\xi,X+\varepsilon\delta X,u)=\mathsf{G}_{X}\big(\frac{\partial H}{\partial X},\delta\xi\big)=\int_{D}[\frac{\partial H}{\partial X}(\xi,X,u)\cdot\delta X]\rho_{0}(a)da$
which implies that
$\frac{\partial H}{\partial\xi}=u(X),\quad\frac{\partial H}{\partial X}=\nabla^{*}u(X)\xi+\frac{1}{2}\lvert u(X)\rvert^{2}(\nabla^{*}X)^{-1}\nabla\log\rho_{0}$
and
$\frac{\partial H}{\partial u}=u\rho-(\rho_{0}\xi)\circ A-\nabla q.$
@@
(III) We consider the following general version of (37):
$\dot{w}+\nabla^{*}u(X)w=f(X),\quad w(0,\cdot)=w_{0}$
Assuming sufficient regularity on $X$, applying Duhamel’s formula we have the following Lagrangian formulae for solution $w(t,\cdot)$:
$w(t,x)=(\nabla^{*}X_{t})^{-1}\bigg(w_{0}(x)+\int_{0}^{t}(\nabla^{*}X_{\tau})f(t,X_{\tau}(x))d\tau\bigg)$
@@
And we are ready to conclude:
$u=\mathbf{P}_{\rho}(\rho^{-1}\eta)=\mathbf{P}_{\rho}\bigg[\nabla^{*}A_{t}\bigg(\xi_{0}-\frac{\nabla\log\rho_{0}}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\bigg)\circ A_{t}\bigg]$ (42)
Combining (38) (42) and $\rho=\rho_{0}\circ A$, we recover the system (31). ∎
@@ 3.2 Vorticity Formulation
We now move on to derive a stream-vorticity formulation of (1). We will state a more general functional analytic framework and treat IIE as a specific example.
Let $\mathcal{V}$ be a Hilbert space with inner product $\langle\cdot,\cdot\rangle_{\mathcal{V}}$. Let $\mathcal{I}:\mathcal{V}\to\mathcal{V}^{*}$ be the Riesz isomorphism between dual pairing $\langle,\rangle_{*}$ of $\mathcal{V},\mathcal{V}^{*}$:
@@
$\langle\varphi,\mathcal{A}\psi\rangle=\langle\varphi,\mathcal{D}^{*}\mathcal{I}\mathcal{D}\psi\rangle$ (44)
Hence we conclude: $\mathcal{A}=\mathcal{D}^{*}\mathcal{I}\mathcal{D}$. While the projection onto the range $\mathsf{Ran}(\mathcal{D})$ is:
@@
In our hydrodynamic setting, assume $d=2$ , let $\mathcal{V}=L^{2}(D;\mathbb{R}^{d})$ be the space of $L^{2}$ vector fields, $\mathcal{F}=H^{1}(D)/\mathbb{R}$, $\mathcal{F}^{*}=H_{0}^{-1}(D)$ be the space of mean-zero $H^{-1}$ distributions. Set $\mathcal{D}=\nabla^{\perp}$, hence $\mathcal{D}^{*}=\nabla\times$. The inertia operator in this case is the following uniformly elliptic operator (due to (A)):
$\mathcal{A}=\nabla\times(\rho\nabla^{\perp})=\mathbf{L}_{\rho}=\Delta_{1/\rho}$
@@
$\mathcal{P}=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho(\cdot))$
Indeed, decomposing $\rho^{-1}w=v+\rho^{-1}\nabla p$ as in proposition 2.3 (i), we have the following
$\mathcal{P}(\rho^{-1}w)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho v)=:\mathcal{K}_{\rho}\nabla\times(\rho v)$
Since $d=2$ and $\nabla\cdot v=0$, consider the stream function $\psi$ of $v$, we find:
@@
$\mathbf{L}_{\rho}:w\in\mathfrak{X}(D)\to\nabla\times(\rho\nabla\times w)$
@@
$\mathcal{F}:=H^{1}(D)/\mathbb{R}$$\mathcal{V}:=L^{2}_{\rho}(D;\mathbb{R}^{2})$$\mathcal{F}^{*}:=H_{0}^{-1}(D)$$\mathcal{V}^{*}\cong L^{2}_{\rho}(D;\mathbb{R}^{2})$$\scriptstyle{\mathcal{D}=\nabla^{\perp}}$$\scriptstyle{\mathcal{I}:\ v\mapsto\rho v}$$\scriptstyle{\mathbf{P}_{\rho}=\mathcal{D}\mathcal{A}^{-1}\mathcal{D}^{*}\mathcal{I}}$$\scriptstyle{\mathcal{A}^{-1}=\Delta_{\rho}^{-1}}$$\scriptstyle{\mathcal{K}_{\rho}=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}}$$\scriptstyle{\mathcal{D}^{*}=\nabla\times}$ (45)
Following the geometric picture above, we now introduce the vorticity formulation. Let $\eta=\rho u+\nabla q$, we therefore have $\mathbf{P}_{\rho}(\rho^{-1}\eta)=u$. Following our above computation, we have yet another closed-form representation of $\mathbf{P}_{\rho}$:
@@
Now we define $\omega=\nabla\times\eta$, then taking the curl of equation (30) we obtain the following vorticity equation:
@@
We note that (47) is algebraically equivalent to system (31)
We now show the conservation of vorticity (Kelvin circulation theorem) in the framework of IIE. In hydrodynamics, conservation laws coming from particle relabeling symmetry (Casimir Invariants in geometry) are known to be the consequence of duality between Lie transport (Lie elements of Lagrangian loops) and dual Lie transport (co-adjoint motion of momentum) ([HMR98][MP25]). Here, for the inhomogeneous fluid, we do not have a full conservation of circulation, due to the commutator forcing term in vortex dynamics. However, the following result holds in $d=2$:
Theorem 3.2
Assume $(u,\rho)$ is a smooth solution of IIE, then for any regular value $\alpha>0$ of $\rho_{0}$, where $E_{\alpha}:=\{\rho_{0}=\alpha\}$ is a smooth loop, let $D_{\alpha}:=\{\rho_{0}<\alpha\}$ be the smooth bounded domain enclosed by $E_{\alpha}$, then we have
$\oint_{X_{t}(E_{\alpha})}u_{t}\cdot d\ell=\oint_{E_{\alpha}}u_{0}\cdot d\ell.$ (48)
@@
Proof. Thanks to the Lagrangian formulae (41), we have for any smooth loop $\gamma$:
@@
$\oint_{E_{\alpha}}u_{t}\cdot d\ell=\alpha^{-1}\oint_{E_{\alpha}}\eta_{t}\cdot d\ell=\alpha^{-1}\oint_{E_{\alpha}}\rho_{0}u_{0}\cdot d\ell=\oint_{E_{\alpha}}u_{0}\cdot d\ell.$
@@
Hence, the proof of (48) is complete. To show (49), simply apply the Stokes theorem. ∎

### v3

@@ 3.1 Hamilton-Pontryagin Principle and Lagrangian-Eulerian Formulation
    and conclude that (33) is equivalent to the system (1) with pressure $p$.
  • (II)
    We compute the partial derivatives $\partial_{\xi}H$, $\partial_{X}H$. Recall the definition of variational derivative:
    $\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}H(\xi+\varepsilon\delta\xi,X,u)=\int_{D}[\frac{\partial H}{\partial\xi}(\xi,X,u)\cdot\delta\xi]\rho_{0}(a)da,$
@@
    $\frac{d}{d\varepsilon}\bigg\lvert_{\varepsilon=0}H(\xi,X+\varepsilon\delta X,u)=\mathsf{G}_{X}\big(\frac{\partial H}{\partial X},\delta X\big)=\int_{D}[\frac{\partial H}{\partial X}(\xi,X,u)\cdot\delta X]\rho_{0}(a)da;$
    which implies the following Hamilton’s ODE
    $\frac{\partial H}{\partial\xi}=u(X),\quad\frac{\partial H}{\partial X}=\nabla^{*}u(X)\xi+\frac{1}{2}\lvert u(X)\rvert^{2}(\nabla^{*}X)^{-1}\nabla\log\rho_{0};$
    with control
    $\frac{\partial H}{\partial u}=u\rho-(\rho_{0}\xi)\circ A-\nabla q=0.$
@@
  • (III)
    We consider the following general version of (31):
    $\dot{w}+\nabla^{*}u(X)w=f(X),\quad w(0,\cdot)=w_{0}.$
    Provided sufficient regularity on $X$, applying Duhamel’s formula we have the following Lagrangian formula for solution $w(t,\cdot)$:
    $w(t,x)=(\nabla^{*}X_{t})^{-1}\bigg(w_{0}(x)+\int_{0}^{t}(\nabla^{*}X_{\tau})f(\tau,X_{\tau}(x))d\tau\bigg)$
@@
    We are now ready to conclude:
    $u=\mathbf{P}_{\rho}(\rho^{-1}\eta)=\mathbf{P}_{\rho}\bigg[\nabla^{*}A_{t}\bigg(\xi_{0}-\frac{\nabla\log\rho_{0}}{2}\int_{0}^{t}\lvert\dot{X}_{\tau}(x)\rvert^{2}d\tau\bigg)\circ A_{t}\bigg].$ (36)
    Combining (32) (36) and $\rho=\rho_{0}\circ A$, we recover the system (25).
@@ 3.2 Vorticity Formulation
We now move on to derive a stream-vorticity formulation of (1). We will state a more general functional analytic framework and treat IIE as a specific example. Here we treat the case that $D$ is simply connected and $D=\mathbb{T}^{2}$ separately.
Let $\mathcal{V}$ be a Hilbert space with inner product $\langle\cdot,\cdot\rangle_{\mathcal{V}}$. Let $\mathcal{I}:\mathcal{V}\to\mathcal{V}^{*}$ be the Riesz isomorphism between dual pairing $\langle\cdot,\cdot\rangle_{*}$ of $\mathcal{V},\mathcal{V}^{*}$:
@@
$\langle\varphi,\mathcal{A}\psi\rangle=\langle\varphi,\mathcal{D}^{*}\mathcal{I}\mathcal{D}\psi\rangle.$ (38)
Hence, we conclude: $\mathcal{A}=\mathcal{D}^{*}\mathcal{I}\mathcal{D}$. While the projection onto the range $\mathsf{Ran}(\mathcal{D})$ is:
@@
In our hydrodynamic setting, assume $d=2$ and $D\subset\mathbb{R}^{2}$ is a simply connected smooth domain. We let $\mathcal{V}=L^{2}(D;\mathbb{R}^{d})$ be the space of $L^{2}$ vector fields, $\mathcal{F}=H^{1}(D)/\mathbb{R}$, $\mathcal{F}^{*}=H_{0}^{-1}(D)$ be the space of $H^{-1}$ distributions. Set $\mathcal{D}=\nabla^{\perp}$, hence $\mathcal{D}^{*}=\nabla\times$. The inertia operator in this case is the following uniformly elliptic operator (due to assumption 2.1):
$\mathcal{A}=\nabla\times(\rho\nabla^{\perp})=:\mathbf{L}_{\rho}=\nabla\cdot(\rho\nabla),$
@@
$\mathcal{P}=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho(\cdot)).$
Here, the inverse $\mathbf{L}_{\rho}^{-1}$ is understood in the sense of Dirichlet boundary condition.
Indeed, decomposing $\rho^{-1}w=v+\rho^{-1}\nabla p$ as in proposition 2.5(I), we have the following identity
$\mathcal{P}(\rho^{-1}w)=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\nabla\times(\rho v)=:\mathcal{K}_{\rho}\nabla\times(\rho v).$ (39)
Here, $\mathcal{K}_{\rho}$ is the modified Biot-Savart operator, which plays the role of the right inverse of the $\rho$-weighted curl operator. (39) is a generalization of the following identity in the homogeneous setting:
$\mathbf{P}v=\nabla^{\perp}\Delta^{-1}\nabla\times v,\quad\forall v\in\mathfrak{X}(D).$
Since $d=2$ and $\nabla\cdot v=0$, let $\psi$ be the stream function of $v$, we find:
@@
$\mathbf{L}_{\rho}:w\in\mathfrak{X}(D)\to\nabla\times(\rho\nabla\times w).$
@@
$\mathcal{F}:=H_{0}^{1}(D)$$\mathcal{V}:=L^{2}_{\rho}(D;\mathbb{R}^{2})$$\mathcal{F}^{*}:=H_{0}^{-1}(D)$$\mathcal{V}^{*}\cong L^{2}_{\rho}(D;\mathbb{R}^{2})$$\scriptstyle{\mathcal{D}=\nabla^{\perp}}$$\scriptstyle{\mathcal{I}:\ v\mapsto\rho v}$$\scriptstyle{\mathbf{P}_{\rho}=\mathcal{D}\mathcal{A}^{-1}\mathcal{D}^{*}\mathcal{I}}$$\scriptstyle{\mathcal{A}^{-1}=\mathbf{L}_{\rho}^{-1}}$$\scriptstyle{\mathcal{K}_{\rho}=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}}$$\scriptstyle{\mathcal{D}^{*}=\nabla\times}$ (40)
Following the geometric picture above, we now introduce the vorticity formulation in this simply-connected setting. Let $\eta=\rho u+\nabla q$, we therefore have $\mathbf{P}_{\rho}(\rho^{-1}\eta)=u$. Following our above computation, we have yet another closed-form representation of $\mathbf{P}_{\rho}$:
@@
Now we define $\omega=\nabla\times\eta$, then taking the curl of (24) we obtain the following vorticity equation:
@@
We note that (42) is algebraically equivalent to system (25). The boundary condition is $u\cdot\mathbf{n}=0$ on $\partial D$, which is compatible with the Dirichlet boundary condition of the Poisson equation
$\mathbf{L}_{\rho}\psi=\omega,\quad\psi\big\rvert_{\partial D}=0.$ (43)
We now show the conservation of vorticity (Kelvin circulation theorem) in the framework of IIE. In hydrodynamics, conservation laws coming from particle relabeling symmetry (Casimir Invariants in geometry) are known to be the consequence of duality between Lie transport (Lie elements of Lagrangian loops) and dual Lie transport (co-adjoint motion of momentum, cf.[24][33]). Here, for the inhomogeneous fluid, we do not have a full conservation of circulation, due to the forcing term in vortex dynamics. However, the following result holds for the case that $D\subset\mathbb{R}^{2}$ is simply connected:
Theorem 3.2 ((Conservation of Circulation)).
Assume $(u,\rho)$ is a smooth solution of IIE. For any regular value $\alpha>0$ of $\rho_{0}$, assume $E_{\alpha}\subset\{\rho_{0}=\alpha\}$ is a smooth loop, let $D_{\alpha}:=\{\rho_{0}<\alpha\}$ be the smooth bounded domain enclosed by $E_{\alpha}$, then we have
$\oint_{X_{t}(E_{\alpha})}u_{t}\cdot d\ell=\oint_{E_{\alpha}}u_{0}\cdot d\ell,$ (44)
@@
Proof of Theorem 3.2.
Thanks to the Lagrangian formula (35), we have for any smooth loop $\gamma$:
@@
$\oint_{X_{t}(E_{\alpha})}u_{t}\cdot d\ell=\alpha^{-1}\oint_{X_{t}(E_{\alpha})}\eta_{t}\cdot d\ell=\alpha^{-1}\oint_{E_{\alpha}}\rho_{0}u_{0}\cdot d\ell=\oint_{E_{\alpha}}u_{0}\cdot d\ell.$
@@
Hence, the proof of (44) is complete. To show (45), simply apply the Stokes theorem to (44) multiplied by $\alpha$.
Remark 3.3 ((Noether Theorem)).
Actually, we may derive the conservation of circulation of IIE as a consequence of Noether’s theorem. By (21), we consider the following action:
$S(\{X_{t}\})=\int_{0}^{T}L(X,\dot{X})dt=\int_{0}^{T}\int_{D}\frac{1}{2}\lvert\dot{X}(t,a)\rvert^{2}\rho_{0}(a)da$
We notice that $S$ is invariant under right action of $\Psi\in\mathsf{Diff}_{\rho_{0}}(D)$, the subgroup of $\mathsf{Diff}(D)$ preserving the initial density $\rho_{0}$. Moreover, by Lemma 2.2, for curve $\Psi_{\varepsilon}\subset\mathsf{Diff}_{\rho_{0}}(D)$ generated by vector field $v$ and starting from identity, we have:
$0=\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\rho_{0}=\frac{d}{d\varepsilon}\bigg\rvert_{\varepsilon=0}\rho_{0}(\Psi_{\varepsilon})\det(\nabla\Psi_{\varepsilon})=\rho_{0}\nabla\cdot v+v\cdot\nabla\rho_{0}=\nabla\cdot(v\rho_{0})$
Hence, $\rho_{0}$-preserving flows are generated by vector field $v$ such that $\nabla\cdot(v\rho_{0})=0$. Now, combined with the $\mathsf{SDiff}(D)$-constraint of the IIE dynamics, we conclude that $\mathscr{G}=\mathsf{SDiff}(D)\cap\mathsf{Diff}_{\rho_{0}}(D)$ is the right gauge group for the IIE action $S$, which is significantly smaller than $\mathsf{SDiff}(D)$ and fails the full Kelvin circulation theorem. In view of the above computation, we have:
$T_{X}\mathscr{G}=\{v\circ X:\nabla\cdot v=v\cdot\nabla\rho_{0}=0\}.$
Differentiating $S$ with right action of $\{\Psi_{\varepsilon}\}\subset\mathscr{G}$, we compute:
$\frac{d}{d\varepsilon}\bigg|_{\varepsilon=0}S(\{X_{t}\circ\Psi_{\varepsilon}\})=\frac{d}{d\varepsilon}\bigg|_{\varepsilon=0}\int_{0}^{T}\int_{D}\frac{1}{2}\lvert\dot{X}(t,\Psi_{\varepsilon}(a))\rvert^{2}\rho_{0}(a)da$
$\int_{0}^{T}\int_{D}\dot{X}(t,a)[\nabla\dot{X}(t,a)v(a)]\rho_{0}(a)dadt$
$\int_{0}^{T}\int_{D}\frac{d}{dt}[\dot{X}(t,a)(\nabla X(t,a)v)]\rho_{0}(a)dadt-\int_{0}^{T}\int_{D}\rho_{0}(a)\ddot{X}(t,a)[\nabla X(t,a)v]da$
$\int_{D}u(T,a)\cdot(X_{T})_{\sharp}(v\rho_{0})da-\int_{D}u_{0}(a)\cdot(v\rho_{0})da-\int_{0}^{T}\int_{D}\nabla p(t,a)(X_{\sharp}v)(a)da$
$\int_{D}u(T,a)\cdot(X_{T})_{\sharp}(v\rho_{0})da-\int_{D}u_{0}(a)\cdot(v\rho_{0})da+\int_{0}^{T}\int_{D}p(t,a)(\nabla\cdot v)\circ X^{-1}(a)da$
$\int_{D}u(T,a)\cdot(X_{T})_{\sharp}(v\rho_{0})da-\int_{D}u_{0}(a)\cdot(v\rho_{0})da$
where we integrate by part, use (12), change the variable and apply the following commutation relation between push-forward and divergence:
$\nabla\cdot(X_{\sharp}v)=(\nabla\cdot v)\circ X^{-1}.$
Hence, we conclude for all $v\in\tilde{\mathfrak{g}}:=T_{e}\mathscr{G}$, the quantity
$\mathscr{I}_{t}:=\int_{D}u(t,a)\cdot[X_{\sharp}(\rho_{0}v)](t,a)da$ (46)
is conserved. Typical example includes the case $d=2$ and $v=F(t,\rho_{0})\nabla^{\perp}\rho_{0}$ for any smooth function $F:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}$. Plugging such $v$ back to (46), we have
$\mathscr{I}_{t}=\int_{D}u(t,a)\nabla^{\perp}\rho(t,a)\rho(t,a)F(\rho(t,a))da,\quad\rho(t,a)=\rho_{0}(X_{t}^{-1}(a)).$
Assume that $\mathsf{Ran}(\rho_{t})=[0,c]$ (here $c$ is independent of $t$ since $\rho_{t}$ is a rearrangement of $\rho_{0}$), by coarea formula we have:
$\mathscr{I}_{t}=\int_{0}^{c}\alpha F(\alpha)\bigg(\int_{\{\rho_{t}=\alpha\}}u(t,a)\cdot\frac{\nabla^{\perp}\rho_{t}}{\lvert\nabla\rho_{t}\rvert}d\mathscr{H}^{1}\bigg)d\alpha=\int_{0}^{c}\alpha F(\alpha)\oint_{\{\rho_{t}=\alpha\}}u(t,a)\cdot d\ell d\alpha$
$\int_{0}^{c}\alpha F(\alpha)\Gamma(t,\alpha)d\alpha,\quad\Gamma(t,\alpha):=\oint_{\{\rho_{t}=\alpha\}}u(t,a)\cdot d\ell.$
Thanks to the conservation of $\mathscr{I}_{t}$ and arbitrariness of $F$, we conclude $\Gamma(t,\alpha)=\Gamma(0,\alpha)$ for all $t$ and $\alpha\in\mathsf{Ran}(\rho_{t})$.
Remark 3.4.
It is a natural question to ask that whether such vorticity formulation would lead the way to the study of finite time blow up of IIE in dimension $2$ and $3$. We remark that the Boussinesq system, which is a fluid equation with active tracers coupled via Buoyancy, is demonstrated in the recent breakthrough [19][20] to blow up in finite time in its local well-posedness class. However, in view of the vortex dynamics, the source term in IIE is much harder to control compared to the Boussinesq case. On the other hand, analysis of lifespan and continuation criteria in IIE based on the vorticity formulation is applicable, see the companion paper [35].
We now discuss the stream-vorticity formulation in torus. Assume $D=\mathbb{T}^{2}$, the key observation is that: For IIE, the conservation of linear momentum reads
$\frac{d}{dt}\int_{D}\rho udx=-\int_{D}u\cdot\nabla(\rho u)dx-\int_{D}\nabla pdx=0.$
Which follows from integrating the first term by part and using mean-free property of gradient fields on torus. However, mean of $u$ is not conserved.
By the above observation, WLOG we assume that $\rho_{0}u_{0}$ is mean zero. Abuse our notation a little bit temporarily, we let
$\eta=\rho u=\nabla^{\perp}\psi+\nabla q.$
Now, the dynamical equation for $\omega$ and $\rho$ keep unchanged:
$\partial_{t}\omega+u\cdot\nabla\omega+\{\frac{1}{2}\lvert u\rvert^{2},\rho\}=0,\quad\partial_{t}\rho+u\cdot\nabla\rho=0,$
where $\{\cdot,\cdot\}$ is the standard Poisson bracket in $2$-d.
The constitutive law becomes different: Taking curl on $\eta$ and making use of $\nabla\cdot u=0$, we have:
$\omega=\Delta\psi,\quad u=\rho^{-1}\nabla^{\perp}\psi+\rho^{-1}\nabla p\quad\Longrightarrow\quad u=\mathbf{P}_{\rho}(\rho^{-1}\nabla^{\perp}\Delta^{-1}\omega).$
Hence, the stream-vorticity formulation of IIE on $\mathbb{T}^{2}$ reads:
$\left\{\begin{aligned} &\partial_{t}\omega+\mathscr{L}_{u}\omega+\{\frac{1}{2}\lvert u\rvert^{2},\rho\}=0;\\ &u=\mathbf{P}_{\rho}(\rho^{-1}\nabla^{\perp}\Delta^{-1}\omega);\\ &\partial_{t}\rho+u\cdot\nabla\rho=0.\end{aligned}\right$. (47)
Supplemented by mean-zero condition of $\rho_{0}u_{0}$. Notice that the above constitutive law is a general law and reduces to the weighted Biot-Savart law in case that $D$ is a simply connected domain:
$\mathbf{P}_{\rho}(\rho^{-1}\nabla^{\perp}\Delta^{-1}\omega)=\mathcal{K}_{\rho}\nabla\times(\rho\rho^{-1}\nabla^{\perp}\Delta^{-1}\omega)=\mathcal{K}_{\rho}\omega=\nabla^{\perp}\mathbf{L}_{\rho}^{-1}\omega.$